OpenStudy (anonymous):

Write a polynomial in standard form with given zeros at -1,-1,and -6

OpenStudy (anonymous):

x-(-1)=? when x=-1?

OpenStudy (anonymous):

Its just when x=-1 but it would be (x+10=0 right?

OpenStudy (anonymous):

(x+1)=0

OpenStudy (anonymous):

using the zero factor property that means:\[(x+1)(x+1)(x+6)=0\] because at leat one of them must equal zero for the whole thing to equal 0 because x+1=0 (solution is -1) x+6=0 (solution is -6)

OpenStudy (anonymous):

can u multiply that out yourself?

OpenStudy (anonymous):

(x+1)(x+1)=x^2+2x+1

OpenStudy (anonymous):

so its actually \[(x^2+2x+1)(x+6)\]

OpenStudy (anonymous):

so the polynomial is \[x^3+8x^2+13x+6\]

OpenStudy (anonymous):

or it could also just be: \[x^2+7x+6\] because (x+1)(x+6)=0